March 3, 2023
Return to the calendar.
RC Meeting Information
Current RC: Iuliia Skorodumina.
No RC meeting today.
Thursday Day (03/02). The beam was down almost the whole shift due to the accelerator issues (less than 1.5 hours of total ABU). The main issue was BLM trips (https://logbooks.jlab.org/entry/4142074). Not a very productive shift. Day Shift summary: https://logbooks.jlab.org/entry/4141934 .
Thursday Swing (03/02). Shift started with no beam as MCC was still troubleshouting the BLM trips. At around 8:30 pm delivery of production beam resumed at 70 uA, which was promising. The issue with BLM trips is not yet resolved, though, and seems to be somehow related to Hall A (see https://logbooks.jlab.org/entry/4142316). PD decided to suspend beam delivery to Hall A till the morning, in order not to hold back beam from the other Halls (https://logbooks.jlab.org/entry/4142311 and https://logbooks.jlab.org/entry/4142312). Swing Shift summary: https://logbooks.jlab.org/entry/4142263 .
Friday Owl (03/03). Without BLM trips from Hall A, we ran smoothly overnight. We started at 70 uA, then went down to 60 uA to observe more beam stability (https://logbooks.jlab.org/entry/4142368 and https://logbooks.jlab.org/entry/4142370). Owl Shift summary: https://logbooks.jlab.org/entry/4142364 . In the morning MCC will resume BLM trips troubleshooting, the options are currently being discussed.
For the current setting (Pm=800 MeV/c), we have collected a total of 6134 mC integrated charge out of the planned 15696 mC. The plan is to continue taking production data for this setting at the highest possible current. Since we do not have any major issues at the moment, we will not have the RC meeting today.
Current Run Plan
The plan is to continue taking production data for the Pm=800 MeV/c at the highest possible current.